### v2

# Towards Human-Level 3D Relative Pose Estimation: Generalizable, Training-Free, with Single Reference

Abstract
Humans can easily deduce the relative pose of a previously unseen object, without labeling or training, given only a single query-reference image pair. This is arguably achieved by incorporating i) 3D/2.5D shape perception from a single image, ii) render-and-compare simulation, and iii) rich semantic cue awareness to furnish (coarse) reference-query correspondence. Motivated by this, we propose a novel 3D generalizable relative pose estimation method by elaborating 3D/2.5D shape perception with a 2.5D shape from an RGB-D reference, fulfilling the render-and-compare paradigm with an off-the-shelf differentiable renderer, and leveraging the semantic cues from a pretrained model like DINOv2. Specifically, our differentiable renderer takes the 2.5D rotatable mesh textured by the RGB and the semantic maps (obtained by DINOv2 from the RGB input), then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, back-propagating the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects, given only a single RGB-D reference, without labeling or training. Extensive experiments on LineMOD, LM-O, and YCB-V show that our training-free method significantly outperforms the state-of-the-art supervised methods, especially under the rigorous Acc@5/10/15 metrics and the challenging cross-dataset settings. The codes are available at https://github.com/ethanygao/training-free_generalizable_relative_pose.
Index Terms:
3D Relative Pose Estimation, Differentiable Renderer, Zero-Shot Unseen Generalization, Single Reference, Label/Training-Free Refinement.
I Introduction
Recent years have witnessed great progress in 3D object pose estimation templates-pose ; nguyen2023nope ; wang2024object ; 3DAHV ; zhao2024dvmnet ; wang2021nemo ; kaushik2024source ; gao2019estimation ; gao2016exploiting ; gao2016symmetric , which estimates the 3D rotation of an object depicted in a query RGB image. As a key to facilitating interaction with real-world objects, 3D object pose estimation attracts increasing attention from various areas including computer vision, virtual/augmented reality, robotics, and human-computer interaction perez2016robot ; robots ; AR . To date, the community shows great interest in generalizable 3D object pose estimation templates-pose ; onepose ; onepose++ ; relpose ; 3DAHV ; zhao2024dvmnet owing to its wide applicability, which focuses on the generalization to previously unseen objects, preferably in a zero-shot manner [^1] We discuss the relatively easier instance- or category-level object pose estimation in the Related Work Sect. II-A and II-B, respectively..
Figure 1: Generalizable object pose estimation with different references, i.e., a CAD model, multiple images, or a single image. CAD models and multi-view references offer rich geometry details, however, scanning the precise CAD model and/or calibrating dense views for multiple images are laborious or even impossible for unseen objects in practice, such as augmented reality. We thus focus on estimating the relative pose w.r.t. a single-view reference following 3DAHV ; zhao2024dvmnet , i.e., the relative pose between a reference-query pair, where we treat the reference pose as canonical without any calibration.
TABLE I: The taxonomy of our method in generalizable pose estimation. For each column, we illustrate the applicability in descending order using the text of Bold, Underlined, and Normal. We also include the Human Intelligence as a reference. The state-of-the-art methods used in our experiments are highlighted with their name.
Method Training Label Reference Query
Modality #Instance Modality #Instance
Human Intelligence training-free label-free RGB single RGB single
zeropose ; zephyr ; lin2023sam6d supervised pose CAD multiple RGB-D single
templates-pose ; osop ; megapose ; ornek2023foundpose ; nguyen2023gigapose supervised pose CAD multiple RGB single
fs6d ; latentfusion supervised pose RGB-D multiple RGB-D single
RelPose++ relpose++ , gen6d ; onepose ; onepose++ ; relpose supervised pose RGB multiple RGB single
LoFTR loftr supervised pose+depth RGB single RGB single
3DAHV 3DAHV , DVMNet zhao2024dvmnet supervised pose RGB single RGB single
ZSP zsp training-free label-free RGB-D single RGB-D multiple
Ours training-free label-free RGB-D single RGB single
Existing generalizable 3D object pose estimation methods can be categorized according to how they exploit the reference information, i.e., using a CAD model, multiple images, or a single image as references, as shown in Fig. 1. Specifically, most existing methods leverage a 3D CAD model megapose ; templates-pose ; osop ; zeropose ; zephyr or multiple images latentfusion ; gen6d ; onepose ; onepose++ ; fs6d ; relpose ; relpose++ for template matching or feature extraction, while the requirement of laborious 3D scanning (for the CAD-based methods) or multiple-image pose labeling (for most multi-image methods) severely limits their applicability.
On the other hand, recent methods propose to reframe the generalizable object pose estimation task as relative pose estimation between a query and a reference image from an unseen object, which is termed as generalizable relative object pose estimation 3DAHV ; zhao2024dvmnet . By treating the reference pose as canonical, estimating the relative pose between the reference-query pair successfully bypasses the laborious 3D scanning (of the CAD reference) or dense views calibration (of the multiple-image reference). However, existing methods rely on a large amount of well-labeled poses between the query-reference pairs to effectively train a neural network, thereby imposing the challenge of acquiring high-quantity training data 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ . Moreover, the generalizability of some network-based methods may be impeded by the training data. Our empirical findings suggest that after pretrained on an external large-scale dataset such as Objaverse objaverse , the current state-of-the-art methods 3DAHV ; zhao2024dvmnet require in-dataset finetuning [^2] The in-dataset finetuning denotes that the finetune set comes from the same dataset with the testing set, while not including the testing objects. before testing on unseen objects within the dataset, which might potentially hinder their cross-dataset generalizability.
In this context, we work towards universally applicable zero-shot 3D generalizable relative pose estimation, where i) the object is agnostic/unseen from a cross-dataset, ii) only a single RGB-D image is available for reference without a 3D CAD model or multi-view images, and iii) the ground-truth (relative) pose label is not available. In other words, we aim to establish a novel 3D generalizable (in terms of both objects and datasets) relative pose estimation method given only one reference and one query image, without labeling or training. This is extremely challenging due to the mixture of incomplete shape information and missing reference-query correspondence, which leads to a severely degraded optimization problem.
Our method is inspired by the fact that humans can easily infer the relative pose under the aforementioned rigorous setting, even with large pose differences or severe occlusions. We hypothesize that such intelligence is accomplished through i) perceiving 3D/2.5D shapes from a single image, ii) conducting render-and-compare simulations via imagination, and iii) understanding rich semantic cues of the object. For example, given two viewpoints of an unseen animal, humans are able to infer the 3D/2.5D shape of that animal, then identify the correspondences of the animal eyes, noses, ears, etc, and finally rotate and render the 3D/2.5D model until its projection matches the other view. Note that the semantic cues have the potential to deal with the (self-) occluded missing parts, thus enhancing the comparison process, e.g., an animal tail can be simply ignored in the render-and-compare simulations if it only appears in one image and is (self-) occluded in the other.
The above analysis motivates us to break down our difficulties and fulfill those three requirements. Concretely, we achieve this by formulating a label/training-free framework through an off-the-shelf differentiable renderer following the render-and-compare paradigm. Our input shape to the differentiable renderer is an RGB- and semantic-textured 2.5D mesh of the reference (avoiding the difficult 3D hallucination of an unseen object). Based on this, we construct a pose refinement framework, where the differentiable renderer takes an initial pose to render projections, then back-propagates the gradients from the projection loss (between the rendered and the query images) to refine the initial pose.
Specifically, our method starts with an RGB-D reference and an RGB query, where their semantic maps can be obtained by leveraging an advanced pretrained model DINOv2 dinov2 with the RGB inputs [^3] Note that our method possesses the potential of using only an RGB reference, please see the discussion in Sect. I-A (Applicability) and Sect. VI (Limitations and Future Works) for more details. Moreover, our method works reasonably well even without the DINOv2 semantic maps on the LineMOD dataset, as illustrated in Table V.. We leverage an easy-to-use differentiable renderer nvdiffrast nvdiffrast , which takes the RGB- and semantic-textured 2.5D mesh of the reference as input, then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The pose refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, which flows the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects from an arbitrary dataset without labeling or training, and naturally generalizes to cross-dataset scenarios.
In summary, we propose a novel 3D generalizable relative pose estimation method, which takes only an RGB-D reference and an RGB query pair, without requiring the ground-truth pose labels or training. We achieve this by formulating a pose refinement framework via an off-the-shelf differentiable renderer under the render-and-compare paradigm. Our method does not involve training a network, which naturally possesses zero-shot generalizability in terms of both unseen objects and datasets. We conducted extensive experiments on LineMOD linemod , LM-O lmo and YCB-V posecnn datasets. The results from our training-free method exhibit significant improvement over the state-of-the-art supervised methods, e.g., for Acc@15 metric on the LineMOD dataset linemod and the YCB-V dataset posecnn , our label- and training-free method outperforms the supervised state-of-the-art results by 29.98% and 14.28%, respectively. Our contributions are three-fold:
  • We propose a novel and simple relative pose estimation method that naturally generalizes to unseen objects, without the need for ground truth pose labels or neural network training.
  • Our method is optimized by an off-the-shelf differentiable renderer and thus training-free. This eliminates the training-data dependency in supervised CNN/ViT-based methods, thus ensuring inherently robust generalizability.
  • Our method employs a render-and-compare framework leveraging 2.5D meshes, thereby avoiding the challenging 3D hallucination of unseen objects. Building upon this, semantic features (such as those from DINOv2) can be integrated as texture maps (via PCA dimensionality reduction) into the render-and-compare process.
I-A Taxonomy and Applicability of Our Method
Taxonomy. The taxonomy of our methods in generalizable pose estimation, in terms of training, labeling, as well as the modality and the number of required instances of the reference and the query images, is illustrated in Table I. Our method falls under the category of label/training-free with a single RGB query and a single RGB-D reference.
Applicability. Among Table I, the proposed method shares the closest setting to the human intelligence on relative pose estimation that is able to generalize to unseen objects from an arbitrary dataset, with only an additional one-time-collection depth map for the reference image.
Our method arguably possesses better applicability compared to the state-of-the-art methods summarized in Table I, as i) unlike supervised in-dataset state-of-the-art methods, our method does not require ground truth (GT) pose annotations for training, where obtaining a large number of GT poses is arguably more challenging than acquiring a single reference depth map. ii) Our method requires only a one-time reference depth annotation, which can be collected and fixed in advance. Moreover, depth sensors are commonly used in our primary application domain, i.e., robotics. We have testified in supplementary material Table S1 that our method can still deliver good estimations with an imprecise depth map, simulating noisy depth sensors. iii) Our training-free method naturally generalizes to unseen objects because it does not involve training a neural network, and thus is training-data independent. In contrast, supervised state-of-the-art methods typically perform less satisfactorily in cross-dataset scenarios, as they suffer from training-data dependency, leading to generalization issues given different training and evaluation datasets.
To further examine the potential of fully distilling human intelligence, we conducted ablation experiments in the supplementary material by substituting the GT depth with predictions from the state-of-the-art Depth Anything v2 yang2024depth . We note that to preserve the object shape, our method requires the metric depth, meaning the estimated depth $z$ and spatial dimensions $x,y$ should share the same unit of measurement (e.g., both in meters). However, as shown in Table S2-S4 in supplementary material, the off-the-shelf metric depth estimator occasionally fails to generalize across different datasets, likely due to an imprecisely recovered depth scale caused by varying camera parameters and/or diverse objects across different training and evaluation datasets. We note that a generalizable metric depth estimator would alleviate this issue, but training a generalizable metric depth estimator is beyond the scope of, and may introduce distractions to, our current focus.
Finally, our method also incorporates the segmentation maps of both query and reference objects as input, which can be obtained by pretrained segmentation models such as SAM sam , FastSAM zhao2023fast and Grounded SAM ren2024grounded . We chose not to delve into these segmentation techniques extensively either, for the same sake of minimizing potential distractions.
II Related work
II-A Instance-level 6D Pose Estimation
Current object pose estimation can be categorized into instance-level, category-level, and generalizable methods based on different problem formulations. For instance-level methods, there are roughly three categories: direct regression-based, correspondence-based, and refinement-based. Direct regression-based methods bb8 ; ssd6d ; posecnn ; yolo6d ; efficientpose predict the object pose directly through a neural network. Correspondence-based methods pvnet ; pvn3d ; cdpn ; dpod ; epos ; gdrnet ; sopose ; zebrapose ; cvpr2023keypoint ; checkerpose estimate the 2D-3D/3D-3D correspondence between the 2D images and 3D object models, followed by PnP solvers Epnp to calculate 6D poses. Additionally, refinement-based methods deepim ; rnnpose ; cir incorporate refinement-based steps to improve the prediction performance. However, instance-level methods are trained on instance-specific data and rely heavily on CAD models to render numerous training data. Consequently, their application is limited to the objects on which they were trained.
II-B Category-level 6D Pose Estimation
In category-level methods, the test instances are not seen during training but belong to known categories. Most methods achieve this by either alignment or directly regressing. Alignment-based methods lee2021category ; nocs ; socs ; spd ; sgpa ; dpdn first propose a Normalized Object Coordinate Space (NOCS) nocs as a canonical representation for all possible object instances within a category. A network is then trained to predict the NOCS maps and align the object point cloud with the NOCS maps using the Umeyama algorithm umeyama1991least to determine the object pose. This method typically constructs the mean shape of specific categories as shape priors using offline categorical object models, and the networks are trained to learn deformation fields from the shape priors to enhance the prediction of NOCS maps. In contrast, directly regressing methods cass ; dualposenet ; fsnet ; vinet ; cps++ avoid the non-differentiable Umeyama algorithm and often focus on geometry-aware feature extraction. For instance, CASS cass contrasts and fuses shape-dependent/pose-dependent features to predict both the object’s pose and size directly. Fs-net fsnet leverages 3D Graph Convolution for latent feature extraction, and designs shape-based and residual-based networks for pose estimation. However, while category-level methods strive to address different instances within the same category, their capacity to predict the poses of objects from entirely new categories remains limited, highlighting the ongoing need to broaden the scope of object pose estimation to encompass unfamiliar objects.
II-C Generalizable 6D Pose Estimation
Generalizable algorithms aim to enhance the generalizability of unseen objects without the need for retraining or finetuning. Methods in this category can be classified as CAD-based templates-pose ; osop ; zeropose ; zephyr ; ornek2023foundpose ; megapose ; nguyen2023gigapose ; caraffa2023object or multi-view reference-based latentfusion ; gen6d ; onepose ; onepose++ ; fs6d ; bundlesdf .
For CAD-based approaches, CAD models are often used as prior knowledge for direct feature matching or template generation. In particular, ZeroPose zeropose performs point feature extraction for both CAD models and observed point clouds, utilizing a hierarchical geometric feature matching network to establish correspondences. Following ZeroPose, SAM-6D lin2023sam6d proposed a two-stage partial-to-partial point matching model to construct dense 3D-3D correspondence effectively. Instead, Template-Pose templates-pose utilizes a CAD model to generate a collection of templates and selects the most similar one for a given query image. Similarly, OSOP osop renders plenty of templates and estimates the 2D-2D correspondence between the best matching template and the query image to solve the object pose. MegaPose megapose proposed a coarse network to classify which rendered image best matches the query image and generate an initial pose. Subsequently, multi-view renderings of the initial pose are produced, and a refiner is trained to predict an updated pose.
Multi-view reference-based methods can be further divided into feature matching-based and template matching-based approaches. For the former, multi-view reference-based feature matching methods mainly aim to establish 2D-3D correspondences between the RGB query image and sparse point cloud reconstructed by reference views or 3D-3D correspondences between the RGB-D query and RGB-D reference images. For instance, FS6D fs6d designed a dense prototype matching framework by extracting and matching dense RGBD prototypes with transformers. After the correspondence is established, Umeyama umeyama1991least algorithms are utilized for pose estimation. OnePose/OnePose++ onepose ; onepose++ apply the Structure from Motion (SfM) method to reconstruct a sparse point cloud of the unseen object using all reference views. They then employ an attention-based network to predict the correspondence between 2D pixels and the reconstructed point clouds to estimate the object pose. For the latter, multi-view references can be reviewed as templates for retrieval when plenty of views exist, or used to reconstruct the 3D object models for template rendering, similar to the CAD-based methods. Gen6D gen6d selects the closest reference view for the query image, and then refines the pose through the 3D feature volume constructed from both the reference and query images. Notably, Gen6D requires more than 200 reference images for initial pose selection. On the contrary, LatentFusion latentfusion reconstructs a latent 3D representation of an object to present an end-to-end differentiable reconstruction and rendering pipeline, and then estimates the pose through gradients update. Since a 3D object representation can be reconstructed utilizing the multi-view information, FoundationPose wen2023foundationpose proposed a unified framework to support both CAD-based and multi-view supported setups. When no CAD model is available, they leverage multi-view references to build a neural implicit representation, which is then used for render-and-compare.
II-D Generalizable Relative Pose Estimation
Recent methods 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ highlight the importance of formulating object pose estimation as a relative pose estimation problem. Specifically, 3DAHV ; zhao2024dvmnet address situations where only a single-view reference image is available. 3DAHV evidence that some state-of-the-art feature matching approaches sarlin2020superglue ; loftr ; zsp fail to generate reliable correspondence between the reference-query pair, while energy-based methods relpose ; relpose++ struggles to capture 3D information. Instead, 3DAHV 3DAHV introduces a framework of hypothesis and verification for generating and evaluating multiple pose hypotheses. Based on that, DVMNet zhao2024dvmnet directly lifts the 2D image features to 3D voxel information in a hypothesis-free way, computing the relative pose estimation in an end-to-end fashion by aligning the 3D voxels.
Figure 2: The overview of the proposed method. Given an RGB-D reference and an RGB query, we extract the semantic maps from a pretrained DINOv2 model dinov2 for both reference and query. Then, the reference 2.5D front-surface mesh is reconstructed by the depth input without hallucination, which is subsequently texture-mapped by its RGB and semantic images. By leveraging a differentiable renderer nvdiffrast , we generate the rendered RGB and semantic maps using the textured 2.5D reference mesh under a novel view/pose. Finally, the rendered RGB and semantic maps are compared to their query counterparts, producing losses and back-propagating the gradients through the differentiable renderer to refine the relative pose.
III Method
Following the render-and-compare paradigm, current generalizable pose estimation methods often rely on rotatable 3D CAD models or well-calibrated multi-view images, imposing challenges to acquire the 3D CAD models or expensive pose calibration, especially for previously unseen objects. We instead focus on the generalizable relative pose estimation formulated in the pioneering works 3DAHV ; zhao2024dvmnet , which aims to estimate the relative pose between a reference-query pair, using only a single reference with an arbitrary pose as canonical (without calibration). Our method differs from 3DAHV ; zhao2024dvmnet in not requiring labeled relative pose to train an estimation network.
III-A Overview
Taking an RGB query and an RGB-D reference as input, our method establishes a refinement optimization under the render-and-compare framework, by leveraging a 2.5D (i.e., RGB-D) shape of the reference, a pair of semantic maps for both the query and the reference acquired by a pretrained DINOv2 model dinov2 along with the corresponding RGB maps, and a differentiable renderer to backpropagate the gradients. Note that the 2.5D shape is exploited due to the inherent difficulty of accurately hallucinating the 3D shape of unseen objects when relying solely on a single RGB-D image. This challenge further complicates the task of relative pose estimation, as the hallucinated 3D shape must align precisely with the query to achieve a successful estimation.
Formally, by using subscript to denote query or reference, our method starts with an RGB pair $I_{r}$ and $I_{q}$ for both reference and query, as well as a depth map $D_{r}$ for the reference. We proposed to estimate the relative pose between $I_{r}$ and $I_{q}$, assisted by $D_{r}$. To this end, we first infer the semantic maps $S_{r}$ and $S_{q}$ from $I_{r}$ and $I_{q}$ exploiting a pretrained DINOv2 model dinov2 . Then, we construct a 2.5D mesh model $M_{r}$ for the reference object based on $D_{r}$, to formulate an RGB and semantic maps textured 2.5D mesh $\mathcal{M}_{r}=\{M_{r},I_{r},S_{r}\}$. Subsequently, the textured 2.5D reference mesh $\mathcal{M}_{r}$ is rotated with an (arbitrary) initial pose $P$ by a differentiable renderer nvdiffrast to generate novel $I_{r}(P)$ and $S_{r}(P)$. Finally, the generated $I_{r}(P)$ and $S_{r}(P)$ are compared with the query $I_{q}$ and $S_{q}$, producing a refinement loss and consequently back-propagate gradients to $P$ through the differentiable renderer. Our method operates the render-and-compare procedure in a self-supervised and network-free manner, without labeling or training.
The overview of the proposed method is illustrated in Fig. 2. We detail the comprising elements of our method in the following sections, i.e., semantic map estimation in Sect. III-B, textured 2.5D mesh reconstruction in Sect. III-C, and label/training-free refinement via differentiable renderer in Sect. III-D.
III-B Semantic Map Estimation
In order to estimate the relative pose, human intelligence may unconsciously infer the semantics of the reference-query pair. Subsequently, coarse correspondence can be established with those semantics, resulting in three-fold benefits: i) it helps to filter out the large non-overlapped part under a substantial pose difference, ii) alleviates the influence of occlusions, and iii) eases the degraded optimization of the relative pose estimation.
Benefit from the rapid development of large pretrained models, an elegant off-the-shelf semantic feature extractor is available as DINO/DINOv2 dino ; dinov2 , which shows great zero-shot generalizability to diverse (even texture-less) objects (see Fig. 3 for some examples). We thus incorporate the off-the-shelf DINOv2 model dinov2 to acquire the rich semantics of the input unseen objects.
Specifically, we utilize DINOv2 dinov2 as the semantic feature extractor $\mathbf{\Phi(x)}$, which takes an RGB image $I$ to produce a set of semantic features $F\in\mathbb{R}^{w\times h\times d}$. In order to texture $F$ to the 2.5D model and facilitate the novel pose rendering, we use the principal component analysis (PCA) to reduce the dimension of $F$ from $d$ to 3, obtaining a semantic map $S$:
$S=\texttt{PCA}(\mathbf{\Phi}(I)),\quad\texttt{PCA}:\mathbb{R}^{w\times h\times d}\rightarrow\mathbb{R}^{w\times h\times 3}.$ (1)
By feeding Eq. (1) with $I_{q}$ and $I_{r}$, we can obtain the semantic maps for the query and the reference, $S_{q}$ and $S_{r}$, respectively. To ensure the semantic consistency between reference and query images, we first calculate and fix the PCA transformation using the reference image, then apply this transformation to all the query images.
Figure 3: Illustration of the semantic maps estimated by DINOv2 dinov2 , demonstrating promising zero-shot performance even for texture-less objects.
III-C Textured 2.5D Mesh Reconstruction
In this section, we reconstruct a rotatable 2.5D model of the reference given its depth map $D_{r}$, which is subsequently used to generate novel renderings through the differentiable renderer. Note that our design avoids the challenging 3D hallucination of an unseen object from the depth map, as the hallucinated 3D shape must consistently align with the query for relative pose estimation.
Specifically, given the depth map $D_{r}$ of the reference, we lift the coordinates of the image plane into the 3D space and obtain the front surface 2.5D point clouds $X_{r}\in\mathbb{R}^{N\times 3}$. We then reconstruct the corresponding 2.5D mesh $M_{r}$ from $X_{r}$, to facilitate the rasterization in the renderer. Since the $xy$ coordinates of $X_{r}$ are sampled regularly from the 2D grids, reconstructing $M_{r}$ from $X_{r}$ can be easily achieved by the Delaunay triangulations lee1980two . Finally, we texture $M_{r}$ with both color and semantic maps, obtaining $\mathcal{M}_{r}=\texttt{TextMap}(M_{r},I_{r},S_{r})$ for rendering under novel poses.
Note that as discussed in Sect. I-A (Applicability), our method possesses the potential of using only an RGB reference and estimating an imprecise depth map exploiting an off-the-shelf generalizable depth estimator. Good estimation is validated in Sect. S2 of the supplementary material given an imprecise and noisy depth. We leave training a generalizable depth estimator as our future work to avoid possible distractions in this paper.
III-D Label/Training-Free Refinement via Differentiable Renderer
Our last module of label/training-free refinement is constructed by a differentiable renderer, which takes the textured 2.5D reference mesh $\mathcal{M}_{r}$ and a pose $P$ as input, then renders a novel RGB image and a novel semantic map under the view $P$. By implementing the pose $P$ as a random variable, the render-and-compare/reprojection loss can be back-propagated directly to $P$, ensuring the label/training-free and zero-shot unseen generalization merits of our proposed method.
Formally, by assuming a perspective camera, we leverage a recent differentiable renderer nvdiffrast nvdiffrast , denoted as $\mathcal{R}$, to generate novel RGB and semantic maps, $I_{r}(P)$ and $S_{r}(P)$, from the textured 2.5D reference mesh $\mathcal{M}_{r}$, an arbitrary pose $P$, and the camera intrinsic KK [^4] The camera intrinsic $K$ can be obtained from the image EXIF information.:
$I_{r}(P),S_{r}(P)=\mathcal{R}(P,\mathcal{M_{\text{r}}},K)$ (2)
Back Surface Culling. As the reconstructed mesh is only 2.5D representing the front surface, it is crucial to conduct the back-surface culling during the rendering to filter out the incorrect back-facing polygons. Specifically, for every triangle of the mesh, we first calculate the dot product of their surface normal and the camera-to-triangle (usually set to $[0,0,1]$) and then discard all triangles whose dot product is greater or equal to 0 backface_culling . Please also see the ablation with and without the back-surface culling in Table V.
Finally, the pose $P$ can be optimized to align the rendered $I_{r}(P)$ and $S_{r}(P)$ with the query $I_{q}$ and $S_{q}$, with the re-projection loss calculated by:
$L(P)=L_{1}\left\{I_{r}(P);I_{q}\right\}+L_{2}\left\{S_{r}(P);S_{q}\right\},$ (3)
where $L(P)$ is the final loss to optimize the pose $P$, and we implement both losses by the multi-scale structural similarity (MS-SSIM) msssim as the following:
$L_{1}=1-\text{ms-ssim}\left\{I_{r}(P);I_{q}\right\},$ (4)
$L_{2}=1-\text{ms-ssim}\left\{S_{r}(P);S_{q}\right\},$ (5)
Equation (3) enables us to optimize $P$ simply by gradient descent.
Initialization. As revealed in the majority of prior arts deepim ; megapose ; templates-pose ; latentfusion , a good initialization significantly boosts the performance of the render-and-compare framework.
To this end, we implement our initialization by evenly sampling candidate poses on a sphere and chasing the best one. Specifically, we first sample $m$ viewpoints (azimuth and elevation angles) uniformly using a Fibonacci lattice gonzalez2010measurement , then uniformly sample $n$ in-plane rotation angles for each viewpoint, producing t=mnt=m*n poses as the initializing candidates. By rendering both RGB and semantic maps of those candidate poses, we are able to calculate the re-projection loss by Eq. (3) (without back-propagation in this phase) and choose the pose with the minimal loss as our initialization $P^{\text{init}}$.
Given the initialized pose $P^{\text{init}}$, we perform $N$ iterations with gradient back-propagation to carry out the label/training-free refinement via the differentiable renderer. Our algorithm is detailed in Algorithm 1.
Algorithm 1 Generalizable Label/Training-Free Refinement
0: Reference RGB and depth $I_{r},D_{r}$; query RGB $I_{q}$; differentiable renderer $\mathcal{R}$; pretrained DINOv2 model $\mathbf{\Phi}$, iteration quota NN, learning rate $\alpha$, camera intrinsic $K$.
[…]
7:
$P^{\text{init}}=\mathop{\arg\min}_{P^{i}\in\mathbf{P}}{L_{1}\left\{I_{r}(P^{i});I_{q}\right\}+L_{2}\left\{S_{r}(P^{i});S_{q}\right\}}$
$\triangleright$ Label/Training-Free Refinement via Diff. Renderer
[…]
9:for $i<N$ do
10:  $I_{r}(P),S_{r}(P)\leftarrow\mathcal{R}(P,\mathcal{M_{\text{r}}},K)$
11:  $L(P)=L_{1}\left\{I_{r}(P);I_{q}\right\}+L_{2}\left\{S_{r}(P);S_{q}\right\}$
12:  $P\leftarrow\texttt{GradientDescent}(L(P),\alpha)$
13:end for
13:$P$
IV Experiments
In this section, we extensively validate our method on benchmark datasets including the LineMOD linemod , YCB-V posecnn , and LineMOD-Occlusion (LM-O) lmo datasets. We detail the experimental setup in the following.
IV-A Experimental Setups
State-of-the-art Methods for Comparison. As shown in Table I, there does not exist a method applying the challenging setting of label/training-free and a single reference-query pair like ours. Therefore we choose the state-of-the-art methods that share the closest experimental setups, which are ZSP zsp , LoFTR loftr , RelPose++ relpose++ , 3DAHV 3DAHV , and DVMNet zhao2024dvmnet . Specifically, for ZSP, though it was originally proposed to process multiple queries, it is able to accept one RGB-D query as input. We report its performance based on the single RGB-D query and single RGB-D reference pair. For LoFTR, we use its pretrained weights released by the authors loftr . The weights of DVMNet, 3DAHV, and RelPose++ are retrained on-demand to achieve their best performance (for the details, see the following Benchmark Experiments, and the table captions of Table II, Table III and Table IV).
Figure 4: Histograms of the geodesic distance between the sampled reference-query pairs. The in-plane rotation is included in calculating the histograms.
Datasets. The experiments are carried out on three benchmark object pose estimation datasets, i.e., LineMOD dataset linemod comprises 13 real objects, each depicting a single low-textured object on varying lighting conditions with approximately 1,200 images. LineMOD-Occlusion (LM-O) lmo consists of 1,214 images of the 8 occluded objects, extracted from the LineMOD dataset, the average visible fraction of objects in LM-O is 79.45%. YCB-V posecnn encompasses over 110,000 real images featuring 21 objects characterized by severe occlusion and clutter, it exhibits an average visible object fraction of 87.15%.
Evaluation Metric. Following zhao2024dvmnet ; 3DAHV , we report mean angular error across sampled reference-query pairs. We also evaluate on important metrics of Acc@5/10/15/30, i.e., the percentage of the predictions that are within 5/10/15/30, which can be more rigorous (e.g., Acc@5) and better characterize the performance. The degree of the pose difference between the ground truth $R_{gt}$ and the predictions $\hat{R}$ is calculated by the geodesic distance $D$:
$D=\arccos\left((\text{tr}(\Delta R^{T}_{gt}\Delta\hat{R})-1)/2\right)/\pi$ (6)
Benchmark Experiments. The in-dataset networks of the state-of-the-art DVMNet, 3DAHV, and RelPose++ methods need to be trained on the leave-out subset which comes from the same dataset as the testing subset but does not include the testing objects. For a fair comparison, on the LineMOD dataset, we follow the experiments in DVMNet zhao2024dvmnet and 3DAHV 3DAHV to evaluate 5 objects (i.e., benchvise, camera, cat, driller, duck). For the YCB-V experiments, we design a similar training protocol to enable the comparison with DVMNet, 3DAHV, and RelPose++, where we randomly sample 8 objects (i.e., tuna_fish_can, pudding_box, banana, pitcher_base, mug, power_drill, large_clamp, foam_brick) for evaluation, leaving the remaining 13 objects to train these three methods. Following DVMNet zhao2024dvmnet , we evaluate 3 unseen objects on the LM-O dataset (i.e., cat, driller, and duck). Since the challenging LM-O dataset is typically used solely for evaluation, we directly use the same weights for DVMNet and 3DAHV that were trained in the LineMOD experiments.
Since the results on the rigorous metrics of Acc@5/10 are not reported in the 3DAHV 3DAHV and DVMNet zhao2024dvmnet paper, we thus retrain them using their official codes for the Acc@5/10 evaluation.
Moreover, as a label/training-free method, the performance of our method can be assessed on all the objects of LineMOD, YCB-V, and LM-O datasets, without the need to leave out any training data or leverage any external dataset. We report the performance of our method on the complete LineMOD, YCB-V, and LM-O datasets in Tables S5, S6, and S7 of the supplementary material.
In-dataset and Cross-dataset Evaluation. Beyond the unseen objects generalization, we also test the dataset-level generalization for the network-based methods including the state-of-the-art DVMNet zhao2024dvmnet and 3DAHV 3DAHV , reporting both the in-dataset and the cross-dataset performance. In short, in-dataset and cross-dataset differ in whether the network needs to be finetuned on a subset that comes from the same dataset with the testing set (though not including the testing objects). Therefore, a good cross-dataset performance is desirable, as the network only needs to be (pre-) trained once on a large-scale external dataset without finetuning.
Specifically, for the in-dataset experiments, we follow the exact training protocols of DVMNet zhao2024dvmnet and 3DAHV 3DAHV , which first pretrain on an external large-scale dataset Objaverse objaverse then finetune on a certain dataset (e.g., LineMOD or YCB-V). For cross-dataset experiments, we use the pretrained weights from Objaverse directly without finetuning.
Note that our method, ZSP zsp , and LoFTR loftr do not require a finetuning phase before evaluation, suggesting that our method, ZSP, and LoFTR naturally generalize to an arbitrary dataset [^5] This is achieved by that i) the pose estimation phase of our method, ZSP, and LoFTR are general and do not involve learning a network, and ii) they all use generalizable feature extractors, i.e., DINOv2 or LoFTR.
Reference-Query Pair Generation. We follow DVMNet zhao2024dvmnet and 3DAHV 3DAHV to generate the reference-query pairs with sufficient overlaps for training and testing. Specifically, given a reference rotation $R_{r}$ and a query rotation $R_{q}$, we first convert the rotation matrices $R_{r}$ and $R_{q}$ to Euler angles $(\alpha_{r},\beta_{r},\gamma_{r})$ and $(\alpha_{q},\beta_{q},\gamma_{q})$. Since the in-plane rotation $\gamma$ does not influence the overlaps between the reference and query pair, it is set to 0 and converted back to the rotation matrix, i.e., $\tilde{R}=h(\alpha,\beta,0)$ with $h$ being Euler-angle to rotation matrix transformation. The overlap between the query and the reference is measured by the geodesic distance (i.e., the pose difference in degree) between their in-plane-omitted rotation matrices $\tilde{R_{q}}$ and $\tilde{R_{r}}$ using Eq. (6). Finally, following DVMNet zhao2024dvmnet and 3DAHV 3DAHV , we select the sampled pairs with $\tilde{D}$ less than $90^{\circ}$.
Following DVMNet zhao2024dvmnet and 3DAHV 3DAHV , for each object, we generate 1000 pairs for testing, and 20000 pairs for training DVMNet, 3DAHV, and RelPose++. Fig. 4 illustrates the histograms depicting the statistics of the pairwise pose difference (geodesic distance between rotation matrices $R_{r}$ and $R_{q}$) on the three datasets. All the experiments are carried out on the same testing reference-query pairs.
Implementation Details. For semantic feature extraction, we employ the output tokens from the last layer of the DINOv2 ViT-L model dinov2 . We use nvdiffrast nvdiffrast as our differentiable renderer. We uniformly sample $m=200$ viewpoints and $n=20$ in-plane rotations (resulting in 4000 initialization candidates), the maximal iteration number for differentiable rendering is set to $N=30$. To backpropagate the refinement losses, we use an Adam optimizer adam of 0.01 initial learning rate and decay by a ReduceLROnPlateau scheduler. All the experiments are conducted on a single NVIDIA 4090 GPU.
[…]
Method Settings Error↓ Acc @ $t^{\circ}$ (%) ↑
Training Label RGB-D Mean Err 30 15 10 5
[…]
LoFTR pose+depth no 63.88 23.94 10.80 6.82 2.42
RelPose++ pose no 46.60 42.50 15.80
3DAHV (cross-dataset) pose no 69.24 21.20 5.52 2.52 0.44
3DAHV (in-dataset) pose no 42.77 59.16 25.92 11.36 2.16
DVMNet (cross-dataset) pose no 47.47 36.44 13.14 5.92 1.08
DVMNet (in-dataset) pose no 33.28 55.02 22.38 10.66 2.72
[…]
Figure 5: Qualitative results on LineMOD. This figure shows that our method can handle partially occluded and texture-less objects. We use a 3D bbox to denote poses.
IV-B Experimental Results on the LineMOD Dataset
The results on the LineMOD dataset are illustrated in Table II. We paste the performances of RelPose++ from the 3DAHV paper 3DAHV . We leave the Acc@5/10 performance of RelPose++ blank as those were not reported in 3DAHV and the (pre-) training code of RelPose++ on the external large-scale Objaverse dataset is not available. Table II shows that our label and training-free method significantly outperforms the supervised state-of-the-art DVMNet w.r.t. all the metrics. In addition, the experimental results show that the performance of DVMNet and 3DAHV decreases when facing cross-dataset scenarios. In contrast, without training a network, our approach inherently generalizes across diverse datasets. Especially, our method significantly outperforms DVMNet (in-dataset) for 21.6% and 32.08% w.r.t. the rigorous Acc@5/10. The qualitative results of our method are shown in Fig. 5, and comparisons with different methods are presented in Fig. S3 of the supplementary material. Our results on all the LineMOD objects are detailed in Table S5 of the supplementary material.
IV-C Experimental Results on the YCB-V Dataset
To compare with the state-of-the-art DVMNet zhao2024dvmnet , 3DAHV 3DAHV and RelPose++ relpose++ , we follow the protocols discussed in Sect. IV-A (In-dataset and Cross-dataset Evaluation) to obtain the in-dataset and cross-dataset performance of DVMNet zhao2024dvmnet and 3DAHV 3DAHV , while RelPose++ is trained on the YCB-V dataset only. The performance on the YCB-V dataset is reported in Table III, where our method exhibits a significant improvement of 11.02% and 17.83% w.r.t. the state-of-the-art DVMNet (in-dataset), respectively on the challenging Acc@5/10 metrics. We showcase the qualitative results of our method on the YCB-V dataset in Fig. 6, and those across different methods can be found in Fig. S5 of the supplementary material. Our results on all the YCB-V objects are shown in Table S7 of the supplementary material.
TABLE III: Experimental results on YCB-V. The performance of DVMNet, 3DAHV, and RelPose++ is obtained by training on a leave-out subset of 13 objects. Other parameters/symbols are the same as those in Table II.
Method Settings Error↓ Acc @ $t^{\circ}$ (%) ↑
Training Label RGB-D Mean Err 30 15 10 5
[…]
LoFTR pose+depth no 68.65 29.45 13.56 7.9 3.19
RelPose++ pose no 57.41 23.60 7.13 3.28 0.76
3DAHV (cross-dataset) pose no 66.61 35.06 16.18 8.28 1.50
3DAHV (in-dataset) pose no 69.48 44.54 28.41 16.29 3.59
DVMNet (cross-dataset) pose no 54.12 41.28 17.11 9.35 2.53
DVMNet (in-dataset) pose no 48.88 51.71 27.04 14.03 3.16
[…]
IV-D Experimental Results on the LM-O Dataset
TABLE IV: Experimental results on LM-O. LM-O is typically used solely for testing with only 8 objects under severe occlusions. The results of DVMNet and 3DAHV are tested directly using the weights trained on LineMOD. Since the model weights of RelPose++ used in Table II were not released, we do not compare our method with RelPose++ in this experiment. Other parameters/symbols are the same as those in Table II.
Method Settings Error↓ Acc $t^{\circ}$ (%) ↑
Training Label RGB-D Mean Err 30 15 10 5
[…]
LoFTR pose+depth no 68.15 20.63 9.00 4.87 1.87
3DAHV (cross-dataset) pose no 55.05 32.83 9.47 4.40 0.53
3DAHV (in-dataset) pose no 62.30 40.29 10.57 3.84 0.57
DVMNet (cross-dataset) pose no 51.75 35.52 12.94 5.30 1.33
DVMNet (in-dataset) pose no 48.55 38.62 14.14 7.37 1.87
[…]
Figure 6: Qualitative results on YCB-V. This figure shows that our method can handle partially occluded and texture-less objects. We use a 3D bbox to denote poses.
Finally, we carry out the experiments on the challenging LM-O Dataset with severe occlusions. Following DVMNet zhao2024dvmnet , we conduct the experiments on three unseen objects of the LM-O dataset, i.e., cat, driller, and duck. We note that the LM-O dataset is typically used solely for evaluation. Therefore, the results of DVMNet and 3DAHV are evaluated utilizing the weights finetuned on LineMOD. Nevertheless, since the weights of RelPose++ for the LineMOD dataset have not been released yet and LM-O (with only 8 objects) cannot provide sufficient leave-out data to train RelPose++, we thus do not include RelPose++ for comparison. The results from Table IV demonstrate the promising performance of our method on the severely occluded LM-O dataset. We showcase our performance on the LM-O dataset in Fig. 7, and those across different methods are illustrated in Fig. S4 of the supplementary material. Our results on all the LM-O objects can be found in Table S6 of the supplementary material.
Figure 7: Qualitative results on LM-O. This figure shows that our method can handle severely occluded and texture-less objects. We use a 3D bbox to denote poses.
[…]
V Ablation Analysis
TABLE V: The contributions of the proposed comprising elements on the LineMOD dataset.
Metrics Mean Err↓ Acc $30^{\circ}$ Acc $15^{\circ}$ Acc $10^{\circ}$ Acc $5^{\circ}$
w/o culling 38.09 67.46 52.32 40.82 23.58
only RGB 36.26 67.42 50.40 37.70 19.62
only semantic 31.31 69.32 50.86 38.80 19.22
Ours 29.93 72.06 54.90 42.74 24.32
TABLE VI: Ablation analysis of different semantic features on the LineMOD dataset.
Metrics Mean Err↓ Acc $30^{\circ}$ Acc $15^{\circ}$ Acc $10^{\circ}$ Acc $5^{\circ}$
RGB 36.26 67.42 50.40 37.70 19.62
LoFTR 54.23 46.62 25.00 15.26 5.24
RGB + LoFTR 39.63 64.30 45.14 32.42 14.80
SD 38.45 57.94 37.28 26.32 12.12
RGB + SD 33.78 65.72 47.56 36.76 19.90
DINOv2 31.31 69.32 50.86 38.80 19.22
RGB + DINOv2 29.93 72.06 54.90 42.74 24.32
Figure 8: Illustration of different semantic features as texture maps.
We carefully investigate the following issues by ablation: i) the contribution of each comprising element of our method, including the back-surface culling, and the usage of RGB or semantic modality in Sect. V-A; ii) the ablations on different semantic features in Sect. V-B; iii) the effects of different initialization strategies in Sect. V-C; iv) the effects of different refinement iterations in Sect. V-D; v) the inference time statistics of our method and comparison with other baselines in Sect. V-E; and vi) the failure cases illustrations from the LM-O dataset in Sect. V-F.
V-A The Contributions of the Proposed Comprising Elements
Despite the simplicity of our method, we are interested in investigating the influences for each of our comprising elements, namely the back-surface culling, and the usage of RGB or semantic modality. We perform those ablations on the LineMOD, and the results are reported in Table V.
As expected, removing each of our comprising elements results in a decreased performance, because all of them are exploited with clear motivations. Nonetheless, the encouraging observation is that our method is able to deliver promising results using only the RGB modality without the semantic map, which already outperforms the state-of-the-art DVMNet (in-dataset) zhao2024dvmnet in Table II across the rigorous Acc $5^{\circ},10^{\circ},15^{\circ}$, and $30^{\circ}$. This further illustrates the good applicability of our method when the pretrained DINOv2 model is not available.
V-B Ablation on Semantic features
To further investigate the performance incorporating alternative semantic feature representations, we tested semantic features from LoFTR loftr and Stable Diffusion (SD) stablediffusion . Table VI shows that i) LoFTR and SD features are inferior to DINOv2 for this task; ii) RGB-only performance surpasses the results based solely on LoFTR or SD features; iii) complementing LoFTR, SD, or DINOv2 with RGB improves the final performance.
Table VI reveals a significant performance gap among DINOv2, SD, and LoFTR features. To further investigate this, we visualize these features in Fig. 8, where the last three columns are dimensionality-reduced texture maps using PCA. Figure 8 shows that the DINOv2 feature maps best characterize the semantic cues, well complementing the RGB appearance, thus explaining its superior performance.
V-C Effects of Different Initialization Strategies
The pose estimation performance under the render-and-compare paradigm is largely affected by the initialization deepim ; cir ; rnnpose ; megapose ; templates-pose ; latentfusion . In the following, we investigate different initializations including: i) random initialization, where we randomly sample candidate poses and choose the best one; and ii) uniform initialization, where the candidate poses are uniformly sampled from a Fibonacci lattice with in-plane rotations gonzalez2010measurement , as detailed in Sect. III-D (Initialization). Table VII shows the performance of different initialization strategies using the LineMOD dataset, which demonstrates that the uniform initialization outperforms the random initialization.
Moreover, we also perform extensive ablations on the sampling densities using uniform initialization. As shown in Fig. 9, the performance boosts from 0 to 4000 samples, marginally improves from 4000 to 16000, and saturates after 16000 samples. In our experiments, we choose uniform initialization with 4000 samples to balance the performance and the efficiency.
TABLE VII: Effects of different initialization strategies using the LineMOD dataset.
Initial Strategy Sampling Numbers Error↓ Acc @ $t^{\circ}$ (%) ↑
Mean Err 30 15 10 5
Random Init. 4000 31.73 70.90 52.66 40.88 23.08
Uniform Init. 4000 29.93 72.06 54.90 42.74 24.32
Figure 9: Ablation analysis on different number of uniformly initialized samples. We tested sampling numbers of 500, 1000, 2000, 4000, 10000, 16000, 32000 on Mean Error and Acc @ $10^{\circ}$.
Figure 10: Performance trend w.r.t. number of refining iterations. We tested iteration numbers of (5, 10, 20, 30, 50, 80, 100, 120, 150) on Mean Error and Acc @ $10^{\circ}$.
V-D Effects of Different Refinement Iterations.
Figure 10 illustrates the impact of the iteration numbers for our label/training-free refinement using the LineMOD dataset, which demonstrates that the performance boosts from 0 to 30 iterations, marginally improves from 30 to 120, and saturates after 120 iterations. We thus set the iteration number to 30 in our main experiments to achieve a balance between performance and efficiency.
TABLE VIII: Inference time statistics of our method on LineMOD.
Semantic Fea. Extraction Pose Initialization Refinement Total
0.11s 1.18s 1.02s 2.52s
TABLE IX: Inference time comparison on LineMOD.
Method ZSP RelPose++ LoFTR 3DAHV DMVNet Ours
Time 1.72s 0.69s 0.30s 0.04s 0.04s 2.52s
V-E Analysis on the Inference Time
We collect the inference time per reference-query pair, averaged across the LineMOD datasets on a single 4090 GPU. We report the runtime for each stage of our method in Table VIII. Note that the initialization is efficient with much more candidate samples than the refinement, because those initializing candidate samples can be evaluated in parallel without backpropagation. Table VIII demonstrates the efficiency of our method with a per-pair runtime of 2.52 seconds in total.
We also present comparisons of our approach with the baseline methods, in terms of the inference time, in Table IX. While our method is slower in inference than the state-of-the-art feedforward models, our render-and-compare paradigm is training-free. This eliminates the numerous training hours required by the state-of-the-art feedforward methods, and ensures inherent generalization to unseen objects (i.e., training-free brings desirable training-data independence).
V-F Illustrations of the Failure Cases
We show our failure cases on the LM-O dataset in Fig. 11, where there do not exist sufficient overlaps between the query and the reference. We note such an extremely degraded case as our limitation and discuss it in Sect. VI (Limitations and Future Works).
Figure 11: Failure cases of our method on the LM-O dataset, where there do not exist sufficient overlaps between the query and the reference due to severe occlusions.
VI Discussions and Conclusions
Limitations and Future Works. Our method has the following two limitations. Firstly, our method necessitates the depth information of the reference object as an input. To acquire the metric depth of the reference image, we evaluated a state-of-the-art monocular depth estimation algorithm yang2024depth . The results and discussions in Sect. S3 of the supplementary shows that the metric depth estimation occasionally fails to generalize across different datasets, primarily stemming from the inherent metric scale ambiguities under varying camera parameters and diverse objects. This suggests that the current limitations are likely to be overcome once a generalizable metric depth estimator becomes available. Despite this, we note that depth sensors are commonly used in our primary application domain, i.e., robotics. Our empirical results, presented in Table S1 of the supplementary materials, demonstrate that our method remains robust with imprecise depth obtained by a noisy depth sensor (simulated by adding noise to the ground-truth depth).
Secondly, our method is likely to fail in the severely degraded scenario where there do not exist adequate overlaps between the query and the reference (possibly caused by occlusions, e.g., Fig. 11). Future research with simultaneous render-and-compare and object completion (with minimal inconsistent hallucination) is a promising direction to explore.
We also note an additional future direction about adaptively determining the loss weights of the RGB pair and the semantic pair in Eq. (3) (preferably adapting in each refinement step), though we empirically showed that simply using equal weights (i.e., both set to 1) leads to promising results.
Conclusions. In this paper, we addressed the challenging generalizable relative pose estimation under a rigorous circumstance with only a single RGB-D reference and single RGB query pair as input, and the pose label is not a priori. We establish our label- and training-free method following the render-and-compare paradigm, by exploiting i) the 2.5D (i.e., RGB-D) rotatable reference mesh, ii) the semantic maps of both query and reference (extracted by a pretrained large vision model DINOv2), and iii) a differentiable renderer to produce and back-propagate losses to refine the relative pose. We carried out extensive experiments on the LineMOD, LM-O, and YCB-V datasets. The results demonstrate that our label/training-free approach surpasses the performance of state-of-the-art supervised methods, particularly excelling under the rigorous Acc@5/10/15 metrics.